OpenStudy (anonymous):

Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15.

OpenStudy (anonymous):

You can write an equation: y = 85 + 15x

OpenStudy (anonymous):

a. Use function notation that models the total price of the bracelet and how that price is based on the number of charms. Explain the reasoning behind your equation.

OpenStudy (anonymous):

Oh, function notation is just replacing 'y' with 'f(x)' y = 85 + 15x What do you get?

OpenStudy (anonymous):

f(x) = 85+15

OpenStudy (anonymous):

Yes, but keep the 'x': f(x) = 85 + 15x

OpenStudy (anonymous):

Where 'x' represents the number of charms.

OpenStudy (anonymous):

b.What would be a reasonable domain for this function based on this scenario? Explain why this is a proper domain

OpenStudy (anonymous):

c.If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250. d. Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make. As they walked down the hall of the mall, Marco saw this sales flyer for Store B on a bench.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you have so far Lexis1615?

OpenStudy (anonymous):

what he gave me :(

jimthompson5910 (jim_thompson5910):

ok try graphing the equation you have

OpenStudy (anonymous):

i did that already lol i need the other parts . cause its 3 questions but they all ahve letters :( u want me to open a new one and start over??

jimthompson5910 (jim_thompson5910):

so you did part c?

OpenStudy (anonymous):

Yes .

jimthompson5910 (jim_thompson5910):

d. Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make. so you plug x = 5 into your function and evaluate

OpenStudy (anonymous):

okay @jim_thompson5910